OpenStudy (anonymous):

Can I please get a direct explanation to this problem? Apples sell for $1.90 per pound, and bananas sell for $0.75 per pound. Troy bought some apples and some bananas. Together they weighed 3.8 pounds, and cost $5.84. How many pounds of apples and how many pounds of bananas did Troy buy? A. 1.5 pounds of apples; 2.3 pounds of bananas B. 2.6 pounds of apples; 1.2 pounds of bananas C. 1.2 pounds of apples; 2.6 pound of bananas D. 1.9 pounds of apples; 1.9 pounds of bananas

OpenStudy (anonymous):

How are you supposed to know how many apples and banana's he bought?

OpenStudy (anonymous):

you don't need to know

OpenStudy (anonymous):

you can create the equation using the prices

OpenStudy (anonymous):

???

OpenStudy (anonymous):

I don't quite understand...

OpenStudy (anonymous):

then there would be a whole bunch of answers

OpenStudy (anonymous):

You would first develop a system of two equations in two variables. x count pounds apples y count pounds bananas x+y=3.8 accounts for pounds. 1.9x+0.75y=5.84 accounts for cost. Use the pounds equation to substitute either for x or for y, into the cost equation, simplify, and see what matches. TRYING WHAT I SAID:----------------------------------------- x + y = 3.8 gives y = 3.8 - x Now substitute this formula for y into the cost equation. 1.9x + .75*(3.8 - x) = 5.84 1.9x + .75 * 3.8 - .75x = 5.84 1.15 + 2.85 = 5.84

OpenStudy (anonymous):

sorry it took so long... it was a lot to type up.....

OpenStudy (anonymous):

correction: the final answer is 1.15x + 2.85 = 5.84

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

actually so c

OpenStudy (anonymous):

Oh... I understand now...

OpenStudy (anonymous):

:D